We start our move up to the Clay tomorrow, so today was mainly spent getting ready. We disconnected and drained the supply lines for the NAS cooling system.
That hose, along with a bunch of power and data cables, was made into a cable and we started wrapping it around the NAS. The NAS will move up to the Aux building between the telescopes tomorrow.
Another detail to worry about is balancing the NAS so we don’t overload the rotator motors. The crew finished a counter-weight today, and Alan and Tyson installed it.
Meanwhile, we also tested the main chiller which keeps the ASM itself cool.
Emilio wrote some software to monitor and control the system. See below for a video of it in action during the above test.
While everybody else has been getting their hands dirty, I’ve been putting our VisAO camera through some tests (mainly focusing). As I complain about regularly, we never get very much time to see what it can do. Here are some reduced images with our coronagraphic occulting spot.
Tonight’s quote: “Hey Laird, how long do you plan to use this instrument? I mean it’s not just for Jared’s PhD right?” (Jason Lewis)